# countnullcountdimennullnullnulldimennullnullnulldimennullnullnulldimennullnullnulldimen

### v1

@@ 3.2 Minimax analysis
Let $\pi_{\max}=\max_{j=1,\dots,m}\pi\left(v^{(j)}\right)$ and let $j^{*}$ be the index attaining the maximum. We construct a new truthful mechanism that has expected revenue not smaller than that of $D$, by adjusting the allocation and payment rules for the vectors $v^{(j)}$ for $j=1,\dots,m$.
One can adjust the allocation rule $z(v^{(j)})$ to $z^{\prime}(v^{(j)})$ to mimic $z(v^{(j^{*})})$. First, we set
[…]
Next, let $t$, with $0\leq t\leq m-1$, be the number of goods with value $x$ that $z(v^{(j^{*})})$ allocates. For $z(v^{(j)})$ we choose the first $t$ goods to be allocated as well; if one of these is the good with value $y$, we choose the $t+1$-th good as well. It then follows that
$\langle z(v^{(j^{*})}),v^{(j^{*})}\rangle=\langle z^{\prime}(v^{(j)}),v^{(j)}\rangle.$ (3.10)
We’ll give a small example of this transformation. Suppose that $j^{*}=4$ $j=2$, $m=8$ and that $z(v^{(4)})=(0,1,1,1,0,0,1,1)$, so that $t=4$ (the third $1$ corresponds to the good with value $y$ being allocated). Then we obtain $z^{\prime}(v^{(2)})=(1,1,1,1,1,0,0,0)$, that is in $z^{\prime}$ the first, third, fourth and fifth good, having value $x$, get allocated, as well as the second good with value $y$.
We set $\pi_{max}$ as the charged price for every vector $v^{(j)}$ for $j=1,\dots,m$. Then the resulting mechanism has revenue not smaller than that of $D$ and it is still truthful because of (3.10). ∎
Using Claim 3.4, we can write
$\frac{1}{m}\sum_{v:C_{y}(v)\leq 1}\pi(v)\mathbb{P}^{m}_{\alpha}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))$ $=\frac{\pi_{0}}{m}\cdot\alpha^{m}+\pi_{1}\cdot\alpha^{m-1}(1-\alpha).$ (3.11)
By individual rationality, we have that (using (2.10) in the final inequality)
$\pi_{0}\leq\langle z(v^{(0)}),v\rangle\leq m\cdot x(\alpha)\leq m(\mu-d/2).$ (3.12)
The transformation in the proof of Claim 3.4 guarantees that the good with value $y$ either gets allocated for all vectors $v^{(j)}$, for $j=1,\dots,m$, or for none of them.
Case 1: $z_{j}(v^{(j)})=0$ for all $j=1,\dots,m$. In this case the good with value $y$ is not allocated for any of the valuation vectors $v^{(j)}$ for $j\geq 1$. By individual rationality this mean that $\pi_{1}\leq\langle z(v^{(j)}),v\rangle\leq(m-1)\cdot x(\alpha)\leq(m-1)(\mu-d/2)$. In combination with (3.12) and using $\alpha(m)\leq 1$ for all $m$, we then have that $\frac{\pi_{0}}{m}\alpha^{m}+\pi_{1}\alpha^{m-1}(1-\alpha)\leq(\mu-d/2)+(m-1)(\mu-d/2)m^{-(m+1)}e^{-m}\leq(\mu-d/2)(1+m^{-m}e^{-m})$. Note that taking $m\rightarrow\infty$ then gives an upper bound of $\mu-d/2$ on (3.11).
Case 2: $z_{j}(v^{(j)})=1$ for all $j=1,\dots,m$. In this case the good with value $y$ is allocated for every of the valuation vectors $v^{(j)}$ for $j\geq 1$. We next make another case distinction, based on whether or not $z$ allocates no, or at least one, good in $v^{(0)}$.
Subcase 2.A: $z$ allocates no good in $v^{(0)}$. This means that $z_{i}(v^{(0)})=0$ for $i=1,\dots,m$. Because of individual rationality, we then must have that $\pi_{0}=0$. Also because of individual rationality, we know that $\pi_{1}\leq(m-1)x(\alpha(m))+y(\alpha(m))=(m-1)(\mu-d/2)+\mu+d/(2(1-\alpha(m)))$. This implies, again using $\alpha(m)\leq 1$, that $\frac{\pi_{0}}{m}\cdot\alpha^{m}+\pi_{1}\cdot\alpha^{m-1}(1-\alpha)\leq(\mu-d/2)\cdot m^{-m}e^{-m}+\mu\cdot m^{-(m+1)}e^{-m}+d/2$. Taking $m\rightarrow\infty$ in the right-hand side gives an upper bound of $d/2$ on (3.11).
Subcase 2.B: $z$ allocates at least one good in $v^{(0)}$. This means that $z_{k}(v^{(0)})=1$ for some $k\in\{1,\dots,m\}$. Without loss of generality, by renumbering the goods if needed, we may assume that $k=1$. We also know, by assumption of Case 2, that $z$ allocates good $1$, with value $y$, in the valuation vector $v^{(1)}$, i.e., $z_{1}(v^{(1)})=1$. The dominant strategy condition defining a truthful mechanism, applied with $v=v^{(1)}$ and $w=v^{(0)}$ tells us that $\langle z(v^{(1)}),v^{(1)}\rangle-\pi_{1}\geq\langle z(v^{(0)}),v^{(1)}\rangle-\pi_{0}$. Roughly speaking, the seller knows that the mechanism allocates the first good both in the bid vectors $v^{(0)}$ and $v^{(1)}$, so we have to make sure that the seller has no incentive to misreport the value it has for the first good (which is $y$ in $v^{(1)}$). Since the mechanism allocates the first good in $v^{(0)}$, we know that $\langle z(v^{(0)}),v^{(1)}\rangle\geq y(\alpha)$. Furthermore, we also know that $\langle z(v^{(1)}),v^{(1)}\rangle\leq(m-1)x(\alpha)+y(\alpha)$ (which is the trivial upper bound on $\langle z(v^{(1)}),v^{(1)}\rangle$ in case $z$ allocates all goods in $v^{(1)}$). Applying these estimates, we get
$(m-1)x(\alpha)+y(\alpha)-\pi_{1}\geq\langle z(v^{(1)}),v^{(1)}\rangle-\pi_{1}\geq\langle z(v^{(0)}),v^{(1)}\rangle-\pi_{0}\geq y(\alpha)-\pi_{0},$
from which it follows that $\pi_{1}\leq\pi_{0}+(m-1)x(\alpha)\leq\pi_{0}+(m-1)(\mu-d/2)\leq(2m-1)(\mu-d/2)$, where the last inequality follows by recalling that $\pi_{0}\leq m(\mu-d/2)$ because of individual rationality. Plugging this upper bound on $\pi_{1}$ into (3.11), gives us (again also using $\alpha(m)\leq 1$)
$\frac{\pi_{0}}{m}\cdot\alpha^{m}+\pi_{1}\cdot\alpha^{m-1}(1-\alpha)\leq(\mu-d/2)+(2m-1)(\mu-d/2)m^{-(m+1)}e^{-m}\leq(\mu-d/2)(1+(2m-1)m^{-m}e^{-m}).$
Note that when $m\rightarrow\infty$, we obtain an upper bound of $\mu-d/2$ on the right hand side of (3.11).
To summarize, when $m\rightarrow\infty$, we obtain in all cases an upper bound of $\mu-d/2$ or $d/2$. Hence,
$\lim_{m\rightarrow\infty}\frac{1}{m}\sum_{v:C_{y}(v)\leq 1}\pi(v)\mathbb{P}^{m}_{\alpha}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))$ $\leq\max\{\mu-d/2,d/2\}.$ (3.13)
Also recall that, following from (3.9),
$\lim_{m\rightarrow\infty}\frac{1}{m}\sum_{k=2}^{m}\sum_{v:C_{y}(v)=k}\pi(v)\mathbb{P}^{m}_{\alpha}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))=0.$ (3.14)
We then may conclude that $\lim_{m\rightarrow\infty}\frac{\mathrm{REV}(D,\mathbb{P}^{m}_{\alpha})}{m}\leq\max\{\mu-d/2,d/2\}$, as the revenue is the sum of the expressions in (3.13) and (3.14); see (3.7). ∎
Theorem 1.1 and 1.3 based on the results in Section 3.1 and 3.2, are proven in Appendix B.
4 Ratio revenue maximization
Inspired by the optimality of the bundling mechanism in the absolute revenue maximization problem, we investigate the asymptotic performance of the optimal deterministic robust bundling mechanism for the ratio and absolute regret objective as well. That is, we determine the optimal bundling mechanism among the set of all bundling mechanisms (instead of all deterministic mechanisms). In order to do this, we use the solution of the absolute revenue maximization problem, combined with some additional arguments. An intuitive way of looking at the result in Theorem 4.1 is that the seller can guarantee a revenue close to $m(\mu-d/2)$, whereas the optimal mechanism can achieve a revenue close to $m\mu$, so that the ratio results in $1-d/(2\mu)$ and the absolute regret in $d/2$. This intuition fails for finite values of $m$, but becomes valid as $m$ grows large. Formalizing this, though, requires some care. The two proofs differ in the analysis of the term $\text{OPT}(m,\mathbb{P})=\sup_{D^{\prime}\in\mathcal{D}_{m}}\mathrm{REV}(D^{\prime},\mathbb{P})$ in the objectives. [^5] In fact, the proofs also work if we allow randomized mechanisms in the OPT term.
Theorem 4.1.
Let $\mu>0$ and $0<d<2\mu$. Then
$\lim_{m\rightarrow\infty}\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}=\lim_{m\rightarrow\infty}\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}=1-\frac{d}{2\mu},$ (4.1)
where the supremum is, roughly speaking, asymptotically attained by the bundling price as in (3.3).
Theorem 4.2.
Let $\mu>0$ and $0<d<2\mu$. Then
$\lim_{m\rightarrow\infty}\left[\inf_{p_{m}}\sup_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}\right]=\frac{d}{2},\,$ (4.2)
where the infimum is, roughly speaking, asymptotically attained by the bundling price as in (3.3).
Appendix A One-sided concentration under given mean and MAD
In this section, we will give the proof of Proposition 3.1. We will rely on the following version of Chebyshev’s inequality [17] in the proof.
Lemma A.1 (following from Chebyshev’s inequality [17]).
@@ Appendix A One-sided concentration under given mean and MAD
$\mathbb{P}(Y\leq(1-\gamma)m\mu)\leq\frac{m\sigma^{2}}{\gamma^{2}m^{2}\mu^{2}}=\frac{\sigma^{2}}{(\gamma\mu)^{2}m}.$
@@
The rough outline of the proof is as follows. We show that, if one chooses a fixed $t=t(\mu,d,\epsilon)$ large enough, then the contribution of the values greater than or equal to $t$ to the mean $\mu$ of any $\mathbb{P}_{i}\in\mathcal{P}(\mu,d)$ is roughly at most $g(t)\cdot\mu+d/2$ with $g(t)$ small if $t$ is large enough. If we condition the distribution $\mathbb{P}_{i}$ on values smaller than $t$, then the resulting distribution has mean roughly $(1-g(t))\mu-d/2$ and finite variance $\sigma^{2}(t)$. We can then use Chebyshev’s bound to argue that we have concentration around $m((1-g(t))\mu-d/2)$. This will give the desired result.
We continue with the formal proof, for which we will use the following proposition.
Lemma A.2.
@@
$\sup_{\mathbb{P}\in\mathcal{P}(\mu,d)}\mathbb{E}_{\mathbb{P}}[X\cdot\mathds{1}_{\geq t}(X)]=\frac{d}{2(t-\mu)}\cdot\mu+\frac{d}{2},$ (A.1)
@@
where the second inequality holds because $d^{\prime}\leq d$. The bound can be attained by choosing the two-point distribution supported on $\{x(\alpha),y(\alpha)\}$ for which $1-\alpha=d/(2(t-\mu))$. The proof of this claim is the same as the reasoning we gave above for the distribution supported on $\{x^{\prime},y^{\prime}\}$.
Now fix an arbitrary $\epsilon\in(0,1-d/2\mu)$ and choose $t=t(\mu,d,\epsilon)$ such that $d/(2(t-\mu))\leq\epsilon$. Because $0<\epsilon<1-d/2\mu$, this will always result in $t\geq\mu+d/2$. For any given $\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}$, let $\mathbb{P}^{m}_{t}(h(X_{1},\dots,X_{m})\in E):=\mathbb{P}^{m}(h(X_{1},\dots,X_{m})\in E|X_{1}<t,\dots,X_{m}<t)$ for any function $h$ and event $E$. Note that by definition of $\mathbb{P}^{m}_{t}$, for a given $c_{m}\geq 0$ it holds that
$\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}\geq c_{m}\right)\geq\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}\geq c_{m}\ \Big|\ X_{i}<t\ \forall i\right)=\mathbb{P}_{t}^{m}\left(\sum_{i=1}^{m}X_{i}^{t}\geq c_{m}\right)$ (A.2)
with $(X_{1},\dots,X_{m})\sim\mathbb{P}^{m}$ and $(X^{t}_{1},\dots,X^{t}_{m})\sim\mathbb{P}^{m}_{t}$. Because of Lemma A.2, and using $\mathbb{P}_{i}(X_{i}<t)\leq 1$, we have
$\mu^{t}_{i}:=\mathbb{E}_{\mathbb{P}_{i}}[X_{i}|X_{i}<t]=\frac{\mu-\mathbb{E}_{\mathbb{P}_{i}}[X_{i}\cdot\mathds{1}_{\geq t}(X_{i})]}{\mathbb{P}_{i}(X_{i}<t)}\geq(1-\epsilon)\mu-\frac{d}{2},$ (A.3)
@@
$(\sigma^{t}_{i})^{2}:=\mathbb{E}_{\mathbb{P}_{i}}\left((X^{t}_{i}-\mu^{t}_{i})^{2}|X_{i}<t\right)\leq\mu^{t}_{i}(t-\mu^{t}_{i}),$ (A.4)
@@
$(\sigma^{t}_{i})^{2}\leq\frac{1}{2}t(t-\frac{1}{2}t)=\frac{1}{4}t^{2}.$ (A.5)
Then Chebyshev’s bound in Lemma A.1 with (A.5) tells us that
$\mathbb{P}_{t}^{m}\left(\sum_{i=1}^{m}X_{i}^{t}>(1-\epsilon)m\mu^{t}_{\textup{min}}\right)\geq 1-\frac{t^{2}}{4(\epsilon\mu^{t}_{\textup{min}})^{2}m},$ (A.6)
Combined with (A.2) and (A.3), this then tells us that
$\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}>(1-\epsilon)m\left((1-\epsilon)\mu-\frac{d}{2}\right)\right)\geq 1-\frac{t^{2}}{4(\epsilon((1-\epsilon)\mu-\frac{d}{2}))^{2}m}.$ (A.7)
As $\mathbb{P}^{m}$ was chosen arbitrarily, this inequality also holds for the infimum over $\mathbb{P}^{m}$. The existence of $f(\mu,d,\epsilon)$ then follows. ∎
Appendix B Proof of Theorems 1.1 and 1.3
The proof of Theorem 1.1 follows by observing that
$\mu-\frac{d}{2}$ $=\lim_{m\rightarrow\infty}\displaystyle\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}$
$\leq\lim_{m\rightarrow\infty}\displaystyle\sup_{D\in\mathcal{D}_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{m}$
$\leq\lim_{m\rightarrow\infty}\displaystyle\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\sup_{D\in\mathcal{D}_{m}}\,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{m}$
$\leq\mu-\frac{d}{2}.$
Here the first equality follows from Lemma 3.2. The first and second inequality are standard. The last inequality follows from the fact that $\max\{\mu-d/2,d/2\}=\mu-d/2$ when $0<d\leq\mu$. The observation in Corollary 1.2 follows from the fact that Lemma 3.2 is also true when $\mathcal{P}(\mu,d)$ is replaced by $\mathcal{P}_{2}(\mu,d)$, in combination with the fact that the proof of Theorem 3.3 works with two-point distributions. The proof of Theorem 1.3 follows in a similar fashion, where the last inequality in the statement of Theorem 1.3 is obtained by observing that $\max\{\mu-d/2,d/2\}=d/2$ when $\mu\leq d<2\mu$.
Appendix C Proof of Theorem 4.1
We will prove Theorem 4.1 by proving the following inequalities:

### v2

@@ 4 Proof of the Minimax Revenue Theorem
Since $A\in\mathcal{A}_{m}$ is truthful, we have due to incentive compatibility that
$\langle z(v^{(n,j)}),v^{(n,j)}\rangle-\pi(v^{(n,j)})\geq\langle z(v^{(n-1,j)}),v^{(n,j)}\rangle-\pi(v^{(n-1,j)}),$
and
$\langle z(v^{(n-1,j)}),v^{(n-1,j)}\rangle-\pi(v^{(n-1,j)})\geq\langle z(v^{(n,j)}),v^{(n-1,j)}\rangle-\pi(v^{(n,j)}).$
Adding the incentive compatibility constraints results in
$\langle z(v^{(n,j)}),v^{(n,j)}\rangle+\langle z(v^{(n-1,j)}),v^{(n-1,j)}\rangle\geq\langle z(v^{(n-1,j)}),v^{(n,j)}\rangle+\langle z(v^{(n,j)}),v^{(n-1,j)}\rangle.$
By rearranging terms we get
$\langle z(v^{(n,j)}),v^{(n,j)}\rangle-\langle z(v^{(n,j)}),v^{(n-1,j)}\rangle\geq\langle z(v^{(n-1,j)}),v^{(n,j)}\rangle-\langle z(v^{(n-1,j)}),v^{(n-1,j)}\rangle,$
or equivalently,
$\langle z(v^{(n,j)}),v^{(n,j)}-v^{(n-1,j)}\rangle\geq\langle z(v^{(n-1,j)}),v^{(n,j)}-v^{(n-1,j)}\rangle.$
This implies that
$\langle z(v^{(n,j)})-z(v^{(n-1,j)}),v^{(n,j)}-v^{(n-1,j)}\rangle\geq 0.$
Since the valuations $v^{(n,j)},v^{(n-1,j)}$ are the same except at location $j$, this results in
$(x_{n}-x_{n-1})\cdot(z_{j}(v^{(n,j)})-z_{j}(v^{(n-1,j)}))\geq 0.$
Since $x_{n}>x_{n-1}$, we have that $z_{j}(v^{(n,j)})\geq z_{j}(v^{(n-1,j)})$. Taking the average (summing over all $j$ and dividing by $m$) gives $\bar{s}_{n}\geq\bar{s}_{n-1}$, which implies $\Delta_{n}\geq 0$. ∎
Next, notice that $\sum_{n=1}^{N}\Delta_{n}=\bar{s}_{N}-\bar{s}_{0}\leq 1-\bar{s}_{0}$. Then, in combination with Lemma 4.1, we get
$\sum_{i=1}^{N}x_{n}w_{n}\Delta_{n}\leq\max\{w_{1}x_{1},\dots,w_{N}x_{N}\}\sum_{n=1}^{N}\Delta_{n}\leq\max\{w_{1}x_{1},\dots,w_{N}x_{N}\}(1-\bar{s}_{0}).$ (4.18)
Taking $N=\lceil\frac{d}{2(\mu-d/2)}\rceil$ and $m\xrightarrow{}\infty$ then gives
$\lim_{m\xrightarrow{}\infty}\frac{1}{m}\pi_{0}+\sum_{n=1}^{N}\bar{\pi}_{n}w_{n}.\leq\lim_{m\xrightarrow{}\infty}x_{0}\bar{s}_{0}+C(m)+\max\{w_{1}x_{1},\dots,w_{N}x_{N}\}(1-\bar{s}_{0})$
$\leq\left(\mu-\frac{d}{2}\right)\bar{s}_{0}+\frac{d}{2N}(1-\bar{s}_{0})\leq\max\left\{\mu-\frac{d}{2},\frac{d}{2N}\right\}=\mu-\frac{d}{2}.$ (4.19)
This completes the proof.
5 Proof of the Maximin Bundling Theorem
In this section we prove the result of Theorem 3.2. First, we derive a one-sided concentration bound for fixed mean and MAD, and then we use this bound to complete the proof.
5.1 Deriving the one-sided concentration bound
The cornerstone of the proof of Theorem 3.2 is Proposition 5.2, which might be of independent interest. Informally, it states that the sum of $m$ i.i.d. random variables with mean $\mu$ and mean absolute deviation (MAD) $d$ will be greater than or equal to $\approx m(\mu-d/2)$ with probability close to one as $m$ grows large. We will use the following one-sided Chebyshev bound in the proof of Proposition 5.2.
Lemma 5.1 (Chebyshev bound).
@@ 5.1 Deriving the one-sided concentration bound
$\mathbb{P}(Y\leq(1-\gamma)m\mu)\leq\frac{m\sigma^{2}}{\gamma^{2}m^{2}\mu^{2}}=\frac{\sigma^{2}}{(\gamma\mu)^{2}m}.$ (5.1)
We now state the one-sided concentration result.
Proposition 5.2 (One-sided concentration bound).
Let $\mu>0$, $0\leq d<2\mu$, and $\epsilon\in(0,\frac{1}{2}(1-\frac{d}{2\mu}))$ be given. Then, for every $m\in\mathbb{N}$, it holds that
$\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}\geq(1-\epsilon)^{2}m\left(\mu-\frac{d}{2(1-\epsilon)}\right)\right)\geq 1-\frac{f(\mu,d)}{m\epsilon^{4}},$ (5.2)
where $f(\mu,d)=\left(\frac{2\mu+d}{2\mu-d}\right)^{2}$ is a function independent of $m$ and $\epsilon$.
@@
The outline of the proof is as follows. We show that, if one chooses a fixed $t=t(\mu,d,\epsilon)$ large enough, then the contribution of the values greater than or equal to $t$ to the mean $\mu$ of any $\mathbb{P}_{i}\in\mathcal{P}(\mu,d)$ is roughly at most $g(t)\cdot\mu+d/2$ with $g(t)$ small if $t$ is large enough. If we condition the distribution $\mathbb{P}_{i}$ on values smaller than $t$, then the resulting distribution has mean $(1-g(t))\mu-d/2$ and finite variance. We can then use Chebyshev’s bound to argue that we have concentration around $m((1-g(t))\mu-d/2)$. This will give the desired result.
We continue with the formal proof, for which we will use the following lemma.
Lemma 5.3.
@@
$\sup_{\mathbb{P}\in\mathcal{P}(\mu,d)}\mathbb{E}_{\mathbb{P}}[X\cdot\mathds{1}_{\geq t}(X)]=\frac{d}{2(t-\mu)}\cdot\mu+\frac{d}{2},$ (5.3)
@@
where the second inequality holds because $d^{\prime}\leq d$. The bound can be attained by choosing the two-point distribution supported on $\{x(\alpha),y(\alpha)\}$ for which $1-\alpha=d/(2(t-\mu))$. The proof of this claim is the same as the reasoning we gave above for the distribution supported on $\{x^{\prime},y^{\prime}\}$. ∎
Now fix an arbitrary $\epsilon\in(0,\frac{1}{2}(1-\frac{d}{2\mu}))$ and choose $t=\mu+d/(2\epsilon)$. Because $\epsilon\in(0,\frac{1}{2}(1-\frac{d}{2\mu}))\subset(0,1)$, this will always result in $t\geq\mu+d/2$. For any given $\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}$, let $\mathbb{P}^{m}_{t}(h(X_{1},\dots,X_{m})\in E):=\mathbb{P}^{m}(h(X_{1},\dots,X_{m})\in E|X_{1}<t,\dots,X_{m}<t)$ for any function $h$ and event $E$. Note that by definition of $\mathbb{P}^{m}_{t}$, for a given $c_{m}\geq 0$ it holds that
$\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}\geq c_{m}\right)\geq\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}\geq c_{m}\ \Big|\ X_{i}<t\ \forall i\right)=\mathbb{P}_{t}^{m}\left(\sum_{i=1}^{m}X_{i}^{t}\geq c_{m}\right)$ (5.4)
with $(X_{1},\dots,X_{m})\sim\mathbb{P}^{m}$ and $(X^{t}_{1},\dots,X^{t}_{m})\sim\mathbb{P}^{m}_{t}$. Because of Lemma 5.3, and using $\mathbb{P}_{i}(X_{i}<t)\leq 1$, we have
$\mu^{t}_{i}:=\mathbb{E}_{\mathbb{P}_{i}}[X_{i}|X_{i}<t]=\frac{\mu-\mathbb{E}_{\mathbb{P}_{i}}[X_{i}\cdot\mathds{1}_{\geq t}(X_{i})]}{\mathbb{P}_{i}(X_{i}<t)}\geq(1-\epsilon)\mu-\frac{d}{2},$ (5.5)
@@
$(\sigma^{t}_{i})^{2}:=\mathbb{E}_{\mathbb{P}_{i}}\left((X^{t}_{i}-\mu^{t}_{i})^{2}|X_{i}<t\right)\leq\mu^{t}_{i}(t-\mu^{t}_{i}),$ (5.6)
@@
$(\sigma^{t}_{i})^{2}\leq\frac{1}{2}t(t-\frac{1}{2}t)=\frac{1}{4}t^{2}.$ (5.7)
Then Chebyshev’s bound in Lemma 5.1 with (5.7) tells us that
$\mathbb{P}_{t}^{m}\left(\sum_{i=1}^{m}X_{i}^{t}>(1-\epsilon)m\mu^{t}_{\textup{min}}\right)\geq 1-\frac{t^{2}}{4(\epsilon\mu^{t}_{\textup{min}})^{2}m},$ (5.8)
Combined with (5.4) and (5.5), this then tells us that
$\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}>(1-\epsilon)m\left((1-\epsilon)\mu-\frac{d}{2}\right)\right)\geq 1-\frac{t^{2}}{\epsilon^{2}(2(1-\epsilon)\mu-d)^{2}m}$
$=1-\frac{(2\mu\epsilon+d)^{2}}{4\epsilon^{4}(2(1-\epsilon)\mu-d)^{2}m}.$ (5.9)
Since $2\mu\epsilon\leq(2\mu-d)/2$ and $2(1-\epsilon)\mu-d\geq(2\mu-d)/2$, (5.1) results in
$\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}>(1-\epsilon)m\left((1-\epsilon)\mu-\frac{d}{2}\right)\right)\geq 1-\frac{(2\mu+d)^{2}}{(2\mu-d)^{2}m\epsilon^{4}}.$ (5.10)
As $\mathbb{P}^{m}$ was chosen arbitrarily, this inequality also holds for the infimum over $\mathbb{P}^{m}$. ∎
5.2 Applying the one-sided concentration result
Recall from (2.2) that $\mathrm{BUND}(p_{m},\mathbb{P}^{m})=p_{m}\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p_{m})$. Using Proposition 5.2, it follows that for every $0<\epsilon<\frac{1}{2}(1-d/2\mu)$, we have
$\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{p_{m}\cdot\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p_{m})}{m}$ $\geq\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{p_{m}^{*}\cdot\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p_{m}^{*})}{m}$
$\geq\frac{p_{m}^{*}}{m}\left(1-\frac{f(\mu,d)}{m\epsilon^{4}}\right).$ (5.11)
Take $\epsilon=\epsilon_{m}=\frac{1}{2}(1-\frac{d}{2\mu})m^{-1/5}$, so that the expression on the right approaches $p_{m}^{*}/m$ as $m\rightarrow\infty$. Note that any $\epsilon$ of the form $\frac{1}{2}(1-\frac{d}{2\mu})m^{-1/(4+\delta)}$ with $\delta>1$ is sufficient.
using that $\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p_{m})\leq 1$. Combining the lower and upper bounds in (5.11) and (5.12), respectively, it follows that
$\mu-\frac{d}{2}=\lim_{m\xrightarrow{}\infty}(1-\epsilon_{m})^{2}\left(\mu-\frac{d}{2(1-\epsilon_{m})}\right)\leq\lim_{m\rightarrow\infty}\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}\leq\mu-\frac{d}{2},$
since $\lim_{m\xrightarrow{}\infty}\epsilon_{m}=\lim_{m\xrightarrow{}\infty}\frac{1}{2}(1-\frac{d}{2\mu})m^{-1/5}=0$. This completes the proof.
6 Unified bundling price in the many-goods limit
We will prove Theorem 3.3 by resolving (3.5) and (3.6) separately while demonstrating that $p^{*}_{m}$ achieves the optimal values.
6.1 Competitive ratio under bundling
We first analyze the competitive ratio by solving (3.5).
Proposition 6.1.
Let $0\leq d<2\mu$. The maximin competitive ratio satisfies
$\lim_{m\rightarrow\infty}\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}=1-\frac{d}{2\mu},$ (6.1)
where the supremum is asymptotically attained by bundling price $p^{*}_{m}$ in (3.3).
Proof.
We will prove (6.1) by proving the following inequalities: